A normal population has a mean of 80.0 and a standard deviation of 14.0.
Compute the probability of a valuebetween 75.0 and 90.0.
Compute the probability of a value75.0 or less.
Compute the probability of a valuebetween 55.0 and 70.0.

Answer by ewatrrr(24785)

 
Hi,
__________________________________________________________________________
HIghly Recommend getting familiar with TI Calculator OR Excel functions/Stat tool
OR the use of various websites for your statistical calculations:Stattrek.com, easycalculation.com ..etc
___________________________________________________________________________
A normal population has a mean of 80.0 and a standard deviation of 14.0.
Compute the probability of a valuebetween 75.0 and 90.0.
P( (90-80)/14 > z >(75-80)/14) =NORMSDIST(10/14) -NORMSDIST(-5/14) = .76 - .36= .40
Compute the probability of a value 75.0 or less.
P(z < (75-80)/14) = NORMSDIST(-5/14) = .36
Compute the probability of a value between 55.0 and 70.0.
P( (55-80)/14 < z <(70-80)/14) =NORMSDIST(-10/14)- NORMSDIST(-25/14) = .24- .04= .20